# WAPITI: A Watermark for Finetuned Open-Source LLMs

### v1

@@ 3.2 Watermark Parameter
$\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{FT}}}(\bm{x})\Delta\bm{\theta}_{\text{Base}}\approx\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{Base}}}(\bm{x})\Delta\bm{\theta}_{\text{Base}}.$ (9)
In this way, we obtain the relationship between the gradient of the fine-tuned model and base models. And we now proceed to derive our target $f_{\bm{\theta}_{\text{FT}}}(\bm{x})$. First, by substituting $g(\bm{x})$ from Eq. (3) into Eq. (2):
$f_{\bm{\theta}_{\text{FT}}^{\dagger}}(\bm{x})=f_{\bm{\theta}_{\text{FT}}}(\bm{x})+\left(\frac{\delta_{\text{FT}}}{\delta_{\text{Base}}}\langle\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{Base}}}(\bm{x}),\Delta\bm{\theta}_{\text{Base}}\rangle+O(\|\Delta\bm{\theta}_{\text{Base}}\|^{2})\right).$ (10)
@@
$f_{\bm{\theta}_{\text{FT}}^{\dagger}}(\bm{x})$ $\approx f_{\bm{\theta}_{\text{FT}}}(\bm{x})+\langle\nabla_{\bm{\theta}}f_{\bm{\theta}_{\text{FT}}}(\bm{x}),\lambda_{\text{FT}}\cdot\Delta\bm{\theta}_{\text{Base}}\rangle+O\left(\|\Delta\bm{\theta}_{\text{Base}}\|^{2}\right),$ (11)
$\approx f_{\bm{\theta}_{\text{FT}}+\lambda_{\text{FT}}\cdot\Delta\bm{\theta}_{\text{Base}}}(\bm{x}).$ (12)
@@
$\bm{\theta}_{\text{FT}}^{\dagger}:=\bm{\theta}_{\text{FT}}+\lambda_{\text{FT}}\cdot\Delta\bm{\theta}_{\text{Base}}.$ (13)
@@
1:base model parameter $\bm{\theta}_{\text{Base}}$, fine-tuned model parameter $\bm{\theta}_{\text{FT}}$, watermark intensity factor $\lambda_{\text{FT}}$
2: watermarked fine-tuned model parameter $\bm{\theta}_{\text{FT}}^{\dagger}$
3:$\bm{\theta}_{\text{Base}}^{\dagger}\leftarrow\text{WatermarkDistillation}(\bm{\theta}_{\text{Base}})$
4:$\Delta\bm{\theta}_{\text{Base}}\leftarrow\bm{\theta}_{\text{Base}}^{\dagger}-\bm{\theta}_{\text{Base}}$
5:$\bm{\theta}_{\text{FT}}^{\dagger}\leftarrow\bm{\theta}_{\text{FT}}+\lambda_{\text{FT}}\cdot\Delta\bm{\theta}_{\text{Base}}$
@@
4 Experiment
4.1 Experimental Setup
@@
Watermark and hyperparameters.  We experiment with two representative decoding-based watermarks, KGW and AAR, with different hyperparameters. To ensure a fair and consistent comparison, we adopt the same watermarking hyperparameters as used by Gu et al. (2024). Specifically, for KGW, we set $k=\{0,1,2\}$, $\gamma=0.25$, and $\delta=\{1,2\}$; and for AAR, we use $k=\{2,3,4\}$. The coefficient $\lambda_{\text{FT}}$ for watermark parameter integration ranges from $[0,4]$.
@@
Dataset and model choices.  To ensure the generalizability of WAPITI, we conduct experiments on two widely used LLM families: Llama-2-7B and Pythia-1.4B, which differ in both architecture and parameter Their popularity in the community further ensures that our experiments reflect real-world utility. We utilize the watermark-distilled base models from Gu et al. (2024).
@@
To test the compatibility of WAPITI with fine-tuned models, we focus on three key fine-tuning capabilities: instruction-following, question answering, and math. We will refer to corresponding fine-tuned models as Llama-chat, Llama-QA, Llama-gsm8k, and Pythia-chat in the experiment results. Detailed information on the fine-tuned model selection can be found in Appendix D. The benchmark datasets used are OpenWebText (Gokaslan & Cohen, 2019), MMLU (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021), respectively.
4.2 Evaluation Metrics
@@
Following the evaluation methods used in Kirchenbauer et al. (2024a), Kuditipudi et al. (2024), and Gu et al. (2024), we evaluate the models on 5,000 samples drawn from the RealNewsLike subset of the C4 dataset (Raffel et al., 2023). The evaluation includes the following metrics:
@@
Watermark detectability.  To assess watermark detectability, we compute the AUROC (Area Under the Receiver Operating Characteristic Curve), which evaluates the ability to distinguish between watermarked and unwatermarked content. The AUROC is calculated using an equal number of human-generated texts and model-generated watermarked content, both truncated to the same length for consistency. Additionally, we compute the median p-values for watermark detection across all generations and use the median p-value as a secondary metric. Lower p-values indicate stronger watermark detectability.
@@
Generation quality.  Generation quality is evaluated using two metrics: perplexity and seq-rep-3 (Sequence Repetition for 3-grams). Perplexity provides an overall assessment of the generated text and is calculated using Llama-2-13B. Seq-rep-3 measures repetition by calculating the proportion of repeated trigrams (Welleck et al., 2019).
@@
Fine-tuning abilities.  To assess whether WAPITI preserves the fine-tuned capabilities of models, we evaluate the performance of WAPITI fine-tuned models on the following benchmarks: i) Question Answering: We use the full MMLU (Hendrycks et al., 2021) dataset to assess the QA ability of models. This dataset contains approximately 14,000 questions from 57 domains. ii) Math: We evaluate the model on the test split of GSM8K (Cobbe et al., 2021), which consists of 1,319 grade-school math word problems designed to assess multi-step reasoning and arithmetic skills.
4.3 Results
Scheme Model Watermark Detectibility Generation Quality
@@ 4.3 Results
None Base Llama $\text{5}\!\cdot\!\text{10}^{-\text{1}}$ 0.50 3.14 0.03
Base Pythia $\text{5}\!\cdot\!\text{10}^{-\text{1}}$ 0.50 10.3 0.04
Table 2: Main results for watermark detectability and generation quality of WAPITI and decoding-based watermarks across different strategies. The displayed results represent the average performance, with an integration coefficient of $\lambda_{\text{FT}}=1$. DECO refers to the original decoding-based watermark used as the baseline.
Watermarking results.  Table 2 presents the results of the watermark strength and generation quality of the WAPITI model. Since multiple hyperparameter sets were tested for each watermarking strategy, the result table displays the average across all hyperparameter sets for each watermark, with the embedded watermark parameter integration coefficient $\lambda_{\text{FT}}$ fixed to 1.0. Detailed results for each hyperparameter, as well as the full set of results for different values of $\lambda_{\text{FT}}$, along with corresponding analysis, can be found in Appendix F.
@@
Of the two watermarks tested, KGW consistently outperforms AAR in watermark transfer, exhibiting higher AUROC scores. This trend is also observed in the watermark-distilled models from Gu et al. (2024), which we partly attribute to the complexity of the AAR scheme, as it combines logits with pseudorandom scores. A more detailed analysis of this difference is provided in Appendix E.1.
Figure 4: Performance of WAPITI models on fine-tuning ability benchmarks are intact after watermarking.
Comparing the performance across different models, the watermark detectability in Pythia models is lower than in Llama models. Analyzing the generations of Pythia models suggests that this difference is largely due to the models’ inherent capabilities. Nevertheless, the parameter integration yields p-values significantly below the baseline of 0.5, indicating that watermarking-related knowledge is still injected to a certain degree.
Fine-tuned ability results.  Figure 4 compares the fine-tuning performance of WAPITI models with the base model and original fine-tuned models. For both QA and Math tasks, WAPITI models show performance nearly identical to the original fine-tuned models for both KGW and AAR watermarking, demonstrating that WAPITI effectively preserves the models’ original capabilities and is fully compatible with fine-tuned models.
@@
4.4 Analysis
In this section, we conduct additional experiments to examine how watermark parameters impact the overall WAPITI models’ watermark detectability and capabilities, providing insights into WAPITI for better utilization and future works.
First, we examine how the norm of watermark parameter integration and the hyperparameters of the watermarking schemes impact detectability and generation quality. We vary the coefficient $\lambda_{\text{FT}}$ within the range $[0,4]$ to test WAPITI models’ median p-values and perplexity. The results show that $\lambda_{\text{FT}}$ regulates the interference between watermarked and model parameters, and detectability strongly correlates with watermark learnability. Full results and analysis are in Appendix E.1.
Second, we evaluate whether WAPITI can defend against fine-tuning attacks by binding fine-tuned capabilities with watermarking. The results show that malicious users would significantly degrade the fine-tuned capabilities of models when attempting to remove the watermark through fine-tuning attacks. Full results can be found in Appendix E.2.
Finally, we analyze the relationship between the fine-tuned and watermarked models at the parameter level using cosine similarity (Ilharco et al., 2023), illustrating how WAPITI remains compatible with fine-tuned models. This analysis also provides strong evidence that watermarked parameters indeed encode knowledge about the watermarking schemes. Full results are detailed in Appendix E.3.
5 Related Work
@@
Text steganography.  Steganography involves embedding information within texts for the purposes of detection or secret communication. Steganography methods can be categorized into edit-based and generative approaches. Edit-based methods include rule-based transformations (Wilson et al., 2014; Wilson & Ker, 2016), synonym-based substitution (Shirali-Shahreza & Shirali-Shahreza, 2008), and neural network-based transformations (Fang et al., 2017; Abdelnabi & Fritz, 2021; Ueoka et al., 2021). On the other hand, generative methods embed information directly during the text generation process (Ziegler et al., 2019; Dai & Cai, 2019).
@@
Text watermarking.  Earlier works in text watermarking typically embedded information through post-processing of texts, closely resembling steganography (Venugopal et al., 2011; Yang et al., 2021). More recent studies have shifted towards decoding-based watermarking, hiding information by perturbing the text during the decoding phase (Kirchenbauer et al., 2024b; Aaronson, 2023; Zhu et al., 2024; Krishna et al., 2023; Kuditipudi et al., 2024; Zhao et al., 2023; Christ et al., 2023; Wu et al., 2024; Liu & Bu, 2024; Giboulot & Teddy, 2024; Lu et al., 2024; Ren et al., 2024; Wang et al., 2024). Different watermarking strategies bring various improvements: Takezawa et al. (2023) enhance logit-perturbation, while Hu et al. (2023); Zhao et al. (2024) optimize sampling strategies. Additionally,  Lee et al. (2024); Li et al. (2023); Yang et al. (2021) explore code watermarking.
@@
Recent advancements have introduced parameter-based watermarking, which embeds watermarks through distillation (Gu et al., 2024). Other studies focus on investigating typical watermarking behaviors (Luo et al., 2024; Singh & Zou, 2023), and some establish robust statistical frameworks for watermarking (Huang et al., 2024; Li et al., 2024). Surveys provide detailed definitions and classifications of text watermarking techniques (Jawahar et al., 2020; Liu et al., 2024; Cai et al., 2024), while benchmarks offer comprehensive evaluations of watermarks (Tu et al., 2024).
@@
Model interventions.  Beyond fine-tuning, researchers have explored parameter-level interventions to modify model behaviors. Key approaches include model patching (Goel et al., 2020; Ilharco et al., 2022; Murty et al., 2022; Sung et al., 2021), parameter editing (Mitchell et al., 2022a; b; Santurkar et al., 2021; Ilharco et al., 2023), and model alignment (Askell et al., 2021; Glaese et al., 2022; Kasirzadeh & Gabriel, 2022). Compared to retraining or fine-tuning, model intervention offers a more efficient way to introduce new capabilities into models.
6 Conclusion

### v2

@@ 3.2. Watermark Parameter
(10) $\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})\Delta\boldsymbol{\theta}_{\text{Base}}\approx\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{Base}}}(\boldsymbol{x})\Delta\boldsymbol{\theta}_{\text{Base}}.$
In this way, we obtain the relationship between the gradient of the fine-tuned model and base models. And we now proceed to derive our target $f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})$. First, by substituting $g(\boldsymbol{x})$ from Eq. (3) into Eq. (2):
(11) $f_{\boldsymbol{\theta}_{\text{FT}}^{\dagger}}(\boldsymbol{x})=f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})+\left(\frac{\delta_{\text{FT}}}{\delta_{\text{Base}}}\langle\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{Base}}}(\boldsymbol{x}),\Delta\boldsymbol{\theta}_{\text{Base}}\rangle+O(\|\Delta\boldsymbol{\theta}_{\text{Base}}\|^{2})\right).$
@@
(12) $f_{\boldsymbol{\theta}_{\text{FT}}^{\dagger}}(\boldsymbol{x})$ $\approx f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x})+\langle\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}_{\text{FT}}}(\boldsymbol{x}),\lambda_{\text{FT}}\cdot\Delta\boldsymbol{\theta}_{\text{Base}}\rangle+O\left(\|\Delta\boldsymbol{\theta}_{\text{Base}}\|^{2}\right),$
(13) $\approx f_{\boldsymbol{\theta}_{\text{FT}}+\lambda_{\text{FT}}\cdot\Delta\boldsymbol{\theta}_{\text{Base}}}(\boldsymbol{x}).$
@@
(14) $\boldsymbol{\theta}_{\text{FT}}^{\dagger}:=\boldsymbol{\theta}_{\text{FT}}+\lambda_{\text{FT}}\cdot\Delta\boldsymbol{\theta}_{\text{Base}}.$
@@
1:base model parameter $\boldsymbol{\theta}_{\text{Base}}$, fine-tuned model parameter $\boldsymbol{\theta}_{\text{FT}}$, watermark intensity factor $\lambda_{\text{FT}}$
2: watermarked fine-tuned model parameter $\boldsymbol{\theta}_{\text{FT}}^{\dagger}$
3:$\boldsymbol{\theta}_{\text{Base}}^{\dagger}\leftarrow\text{WatermarkDistillation}(\boldsymbol{\theta}_{\text{Base}})$
4:$\Delta\boldsymbol{\theta}_{\text{Base}}\leftarrow\boldsymbol{\theta}_{\text{Base}}^{\dagger}-\boldsymbol{\theta}_{\text{Base}}$
5:$\boldsymbol{\theta}_{\text{FT}}^{\dagger}\leftarrow\boldsymbol{\theta}_{\text{FT}}+\lambda_{\text{FT}}\cdot\Delta\boldsymbol{\theta}_{\text{Base}}$
@@
4. Experiment
4.1. Experimental Setup
@@
We experiment with two representative decoding-based watermarks, KGW and AAR, with different hyperparameters. To ensure a fair and consistent comparison, we adopt the same watermarking hyperparameters as used by (Gu et al., 2024). Specifically, for KGW, we set $k=\{0,1,2\}$, $\gamma=0.25$, and $\delta=\{1,2\}$; and for AAR, we use $k=\{2,3,4\}$. The coefficient $\lambda_{\text{FT}}$ for watermark parameter integration ranges from $[0,4]$.
Dataset and model choices.
@@
To ensure the generalizability of WAPITI, we conduct experiments on two widely used LLM families: Llama-2-7B and Pythia-1.4B, which differ in both architecture and parameter Their popularity in the community further ensures that our experiments reflect real-world utility. We utilize the watermark-distilled base models from (Gu et al., 2024).
Evaluation Procedure
We evaluate model generation using samples from the RealNewsLike subset of the C4 dataset (Raffel et al., 2023). Specifically, the evaluation sample size is 5,000, with a 50-token prompt and a sequence length of 200. We use temperature sampling with $t=1$. Building on the approach of  (Gu et al., 2024), we apply deduplication during post-processing to remove repetitive generations, ensuring the validity of the final detectability results.
Evaluation Metrics
@@
To test the compatibility of WAPITI with fine-tuned models, we focus on three key fine-tuning capabilities: instruction-following, question answering, and math. We will refer to corresponding fine-tuned models as Llama-chat, Llama-QA, Llama-gsm8k, and Pythia-chat in the experiment results. Detailed information on the fine-tuned model selection can be found in Appendix D. The benchmark datasets used are OpenWebText (Gokaslan and Cohen, 2019), MMLU (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021), respectively.
4.2. Evaluation Metrics
@@
Following the evaluation methods used in (Kirchenbauer et al., 2024a), (Kuditipudi et al., 2024), and (Gu et al., 2024), we evaluate the models on 5,000 samples drawn from the RealNewsLike subset of the C4 dataset (Raffel et al., 2023). The evaluation includes the following metrics:
Watermark detectability.
@@
To assess watermark detectability, we compute the median p-value and AUROC (Area Under the Receiver Operating Characteristic Curve), which evaluates the ability to distinguish between watermarked and unwatermarked content. The p-value is computed using the z-score method. A lower p-value indicates stronger watermark detectability. The AUROC is calculated using an equal number of human-generated texts and model-generated watermarked content, both truncated to the same length for consistency.
Generation quality.
@@
Generation quality is evaluated using two metrics: perplexity and seq-rep-3 (Sequence Repetition for 3-grams). Perplexity provides an overall assessment of the generated text and is calculated using Llama-2-13B. Seq-rep-3 measures repetition by calculating the proportion of repeated trigrams (Welleck et al., 2019).
Fine-tuning abilities.
@@
To assess whether WAPITI preserves the fine-tuned capabilities of models, we evaluate the performance of WAPITI fine-tuned models on the following benchmarks: i) Question Answering: We use the full MMLU (Hendrycks et al., 2021) dataset to assess the QA ability of models. This dataset contains approximately 14,000 questions from 57 domains. ii) Math: We evaluate the model on the test split of GSM8K (Cobbe et al., 2021), which consists of 1,319 grade-school math word problems designed to assess multi-step reasoning and arithmetic skills.
4.3. Results
Watermark Scheme Model Watermark Detectibility Generation Quality
@@ 4.3. Results
None Base Llama $\text{4.5}\!\cdot\!\text{10}^{-\text{1}}$ 0.48 3.14 0.03
Base Pythia $\text{5.6}\!\cdot\!\text{10}^{-\text{1}}$ 0.49 10.3 0.04
Table 2. Main results for watermark detectability and generation quality of WAPITI and decoding-based watermarks across different strategies. The displayed results represent the average performance, with an integration coefficient of $\lambda_{\text{FT}}=1$. DECO refers to the original decoding-based watermark used as the baseline.
Watermarking results.
Table 2 presents the results of the watermark strength and generation quality of the WAPITI model. Since multiple hyperparameter sets were tested for each watermarking strategy, the result table displays the average across all hyperparameter sets for each watermark, with the embedded watermark parameter integration coefficient $\lambda_{\text{FT}}$ fixed to 1.0.
@@
Of the two watermarks tested, KGW consistently outperforms AAR in watermark transfer, exhibiting higher AUROC scores. This trend is also observed in the watermark-distilled models from (Gu et al., 2024), which we partly attribute to the complexity of the AAR scheme, as it combines logits with pseudorandom scores. A more detailed analysis of this difference is provided in Section 4.5. Comparing the performance across different models, the watermark detectability in Pythia models is lower than in Llama models. Analyzing the generations of Pythia models suggests that this difference is largely due to the models’ inherent capabilities. Nevertheless, the parameter integration yields p-values significantly below the baseline of 0.5, indicating that watermarking-related knowledge is still injected to a certain degree.
Figure 3. Performance of WAPITI models on fine-tuning ability benchmarks are intact after watermarking.
Fine-tuned ability results.
Figure 3 compares the fine-tuning performance of WAPITI models with the base model and original fine-tuned models. For both QA and Math tasks, WAPITI models show performance nearly identical to the original fine-tuned models for both KGW and AAR watermarking, demonstrating that WAPITI effectively preserves the models’ original capabilities and is fully compatible with fine-tuned models.
@@
Robustness to Attacks
WAPITI embeds watermarks through parameter integration, so the watermark parameters must be kept secret, like secret key. Otherwise, malicious users could directly remove the integrated parameters to invalidate the watermark.
Beyond this direct attack, we evaluate WAPITI’s robustness against the classical watermark elimination method: text edits. Detailed experimental setups and robustness analyses are provided in Appendix E.
4.4. Analysis
In this section, we conduct additional experiments to address three core questions: (i) How do watermark parameters affect the watermark detectability and capabilities of WAPITI models? (ii) Does WAPITI offer defense against fine-tuning attacks? (iii) Why is WAPITI compatible with fine-tuned LLMs? These experiments provide insights into WAPITI for improved utilization.
4.5. How will the watermark parameter affect the model performance?
Figure 4. Watermark detectability and output perplexity of the WAPITI model as a function of the watermark integration coefficient $\lambda_{\text{FT}}$ (left and middle). The scatter plot shows the relationship between perplexity and detectability (right).
We evaluated the watermark detectability of the model and generation quality across varying coefficients $\lambda_{\text{FT}}$ for watermark parameter integration. Figure 4 illustrates the watermark detectability (measured by p-value) and perplexity of the WAPITI Llama-math model at different values of $\lambda_{FT}$.
The results in Figure 4(b)(e) show that when the coefficient is in the range [0, 1], watermark strength increases steadily, while perplexity stays below 5.0, indicating minimal impact on the model’s generation capability. Additionally, the gradient of watermark strength in Figures 4(a) and (d) varies with watermarking hyperparameters. For KGW, smaller $k$ reduces the influence of preceding tokens, and larger $\delta$ enhances detectability. Similarly, for AAR, a smaller $k$ also reduces the impact of the previous context. Thus, smaller $k$ and larger $\delta$ make the watermark easier to learn, consistent with (Gu et al., 2024). These trends are evident in Figure 4(a) and (d), where the gradient of watermark strength aligns with watermark learnability.
However, when the coefficient exceeds 1.0, distinct patterns emerge for the two watermarks. For KGW, both watermark detectability and perplexity increase monotonically with the coefficient. AAR shows a parabolic behavior in both strength and perplexity, with their extrema occurring at different values. This divergence suggests that the watermark parameter, while generally applicable, can interfere with the model when $\lambda$ becomes large.
Figure 4(c),(f) show scatter plots of perplexity vs. p-value, illustrating the trade-off between watermark detectability and output quality. Perplexity is constrained within the [0, 20] range to preserve generation quality. KGW exhibits a linear relationship between watermark strength and perplexity, reflecting the expected trade-off. In contrast, the AAR plot shows a more chaotic pattern, with no clear correlation. This difference stems from the watermarking mechanisms: KGW is explicitly decomposable into n-grams, whereas AAR involves both logits and pseudorandom scores, making it more difficult to learn. These results offer insights into the most suitable watermarking strategy for WAPITI.
4.6. Can the watermark vector protect fine-tuned abilities?
Figure 5. The results reveal that fine-tuning attacks degrade both model performance and watermark removal ability, making them an effective defense.
Fine-tuning Attack
Recall the definition in the § 2.1, the Utility of a watermark is defined by the difficulty of removing it without significantly altering the generated content or impairing the inherent capabilities of the model.
Fine-tuning Attack Setup
We use fine-tuned models embedded with each watermark type: KGW with $k=0$, $\gamma=2$, and $\delta=2$. These models are then further fine-tuned on the OpenWebText dataset (Gokaslan and Cohen, 2019) for 1,000 steps. The training configurations are remained the same as (Gu et al., 2024). For each model checkpoint every 200 steps, we generate 200-token completions using prompts of 50 tokens from the C4 RealNewsLike dataset. Then we calculate the model’s fine-tuned capability and its watermark detectability after fine-tuning to show the impact of the fine-tuning attack’s impact.
Defense for Fine-tuning Attack
To defend against fine-tuning attacks, we can bind watermarking to the fine-tuned abilities of the model. In doing so, if malicious users attempt to fine-tune the model to remove the watermark, the fine-tuned capabilities of the model will also be severely compromised, thereby enhancing the robustness of the watermark in fine-tuned models. In this experiment, we select the Llama-Math and Llama-QA, each embedded with the k0-gamma0.25-delta2 watermark. Then we test how watermark strength and fine-tuning performance are affected after additional fine-tuning.
As shown in Figure 5, both models’ watermark detectability and fine-tuning capabilities declined significantly after just 400 steps of fine-tuning attack. This result demonstrates that WAPITI on fine-tuned models is robust to fine-tuning attacks, addressing the vulnerabilities of previous watermark distillation methods.
4.7. How can the watermark vector be compatible with fine-tuned abilities?
Figure 6. The plot shows cosine similarity, indicating clear orthogonality between watermark parameter differences and fine-tuning parameter differences.
Finally, We investigate why the WAPITI is effective across different fine-tuned models by employing a parameter-based approach similar to (Ilharco et al., 2023). We calculate the cosine similarity between the watermark parameter and task vectors (Ilharco et al., 2023), where the task vectors represent the parameter differences between the fine-tuned and base models.
5. Related Work
Text steganography.
@@ 5. Related Work
Steganography involves embedding information within texts for the purposes of detection or secret communication. Steganography methods can be categorized into edit-based and generative approaches. Edit-based methods include rule-based transformations (Wilson et al., 2014; Wilson and Ker, 2016), synonym-based substitution (Shirali-Shahreza and Shirali-Shahreza, 2008), and neural network-based transformations (Fang et al., 2017; Abdelnabi and Fritz, 2021; Ueoka et al., 2021). On the other hand, generative methods embed information directly during the text generation process (Ziegler et al., 2019; Dai and Cai, 2019).
Text watermarking.
@@
Earlier works in text watermarking typically embedded information through post-processing of texts, closely resembling steganography (Venugopal et al., 2011; Yang et al., 2021). More recent studies have shifted towards decoding-based watermarking, hiding information by perturbing the text during the decoding phase (Kirchenbauer et al., 2024b; Aaronson, 2023; Zhu et al., 2024; Krishna et al., 2023; Kuditipudi et al., 2024; Zhao et al., 2023; Christ et al., 2023; Wu et al., 2024; Liu and Bu, 2024; Giboulot and Teddy, 2024; Lu et al., 2024; Ren et al., 2024; Wang et al., 2024). Different watermarking strategies bring various improvements: (Takezawa et al., 2023) enhance logit-perturbation, while (Hu et al., 2023; Zhao et al., 2024) optimize sampling strategies. Additionally,  (Lee et al., 2024; Li et al., 2023; Yang et al., 2021) explore code watermarking.
@@
Recent advancements have introduced parameter-based watermarking, which embeds watermarks through distillation (Gu et al., 2024). Other studies focus on investigating typical watermarking behaviors (Luo et al., 2024; Singh and Zou, 2023), and some establish robust statistical frameworks for watermarking (Huang et al., 2024; Li et al., 2024). Surveys provide detailed definitions and classifications of text watermarking techniques (Jawahar et al., 2020; Liu et al., 2024; Cai et al., 2024), while benchmarks offer comprehensive evaluations of watermarks (Tu et al., 2024).
Model interventions.
@@
Beyond fine-tuning, researchers have explored parameter-level interventions to modify model behaviors. Key approaches include model patching (Goel et al., 2020; Ilharco et al., 2022; Murty et al., 2022; Sung et al., 2021), parameter editing (Mitchell et al., 2022a, b; Santurkar et al., 2021; Ilharco et al., 2023), and model alignment (Askell et al., 2021; Glaese et al., 2022; Kasirzadeh and Gabriel, 2022). Compared to retraining or fine-tuning, model intervention offers a more efficient way to introduce new capabilities into models.
6. Conclusion